# Quantitative stability for quasilinear parabolic equations

### v1

@@ 1.1. Motivation
In this paper we study the stability of a class of quasilinear parabolic partial differential equations, which takes the form
@@
are given functions satisfying appropriate assumptions to be introduced later. Here, ${\mathcal{S}}^{n}_{+}$ represents the set of all nonnegative symmetric $n\times n$ matrices. A key feature of this general setting is that the elliptic operator may exhibit singularity at vanishing gradients.
@@
One of the main questions addressed in this paper concerns the stability of $p$-Laplace type equations under perturbations of the parameter $p$. Assume that $H\equiv 0$. For a fixed parabolic boundary value $g$, it is known that, as $p\to q$ for some $q\in(1,\infty)$, the unique solution $u_{p}$ of the Cauchy-Dirichlet problem (1.2) or (1.3) converges uniformly to the solution $u_{q}$ of the corresponding problem with parameter $q$. In the framework of viscosity solutions, consult [34, Theorem 6.1] and [19, Theorem 2.2.1] for such qualitative stability results on the variational and normalized $p$-Laplace equations of parabolic type. See also [28, 32] for related analysis with different approaches. However, these works do not provide an explicit rate of convergence. Our goal is therefore to further quantify the uniform convergence and to discuss the optimality of the resulting estimates.
On the other hand, when studying the solutions $u$ of singular parabolic equations such as (4.1) and (1.3), it is common to approximate them with regularized equations in the form
@@
where $\varepsilon>0$ is small. Here, we take $p^{\prime}=2$ and $p^{\prime}=p$ to obtain regularized approximations of (4.1) and (1.3), respectively. Under the comparison principle for the limit singular equations, sending $\varepsilon\to 0$ yields the uniform convergence of the solution $u_{\varepsilon}$ of (1.5) to the solution $u$ of the corresponding singular equation. It is then natural to ask about the rate of this convergence in applications. The case with $p^{\prime}=2$ and $p=1$ corresponds to the level-set mean curvature flow equation, for which such a regularization was introduced in [17]. The convergence rate in this case was essentially established by Deckelnick [14], and later proved by Mitake [33] using a different method based on comparison arguments. Addressing convergence rates for the general approximation (1.5) is another main objective of the present paper.
The method employed in this paper builds on the arguments developed by Mitake in [33], as well as earlier works [10, 24], which established general continuous dependence estimates for viscosity solutions of nonlinear parabolic equations. Further related results include extensions to nonlocal PDEs in [25] and to general Neumann-type boundary value problems in [23]. While it may be possible to adapt these methods to study quantitative stability for $p$-Laplace type equations in regimes without vanishing-gradient singularities, precise estimates do not appear to have been explicitly given, to the best of our knowledge. Also, the singular case, such as (1.3) with $1<p<2$, seems unexplored and constitutes one of the focuses of the present work.
A recent paper of Bungert [5] studies the convergence rate of $p$-harmonic functions to infinity-harmonic functions, using partly the variational structure of the $p$-Laplace operator. In contrast, based on the viscosity solution theory, we here provide a purely PDE approach to show the convergence rate for a general class of related parabolic stability problems. Our method develops the techniques in [10, 24, 33] and further extends them to treat convergence rates arising from more general stability and regularization problems.
@@ 1.2. Main results
These assumptions guarantee the existence and uniqueness of $\mathcal{F}$}-solutions to the associated Cauchy-Dirichlet problem. The existence of $k>2$ required in (A1) is related to the possible singularity of the equation at vanishing gradients. A typical example of singular operators we will consider in our later study of the stability is the case of the $p$-Laplacian for $p$ near a fixed $q\in(1,2)$, where $A_{\varepsilon}$ is taken to be
@@ 1.3. Applications
Our general framework in Theorem 1.1 yields stability estimates for a broad class of quasilinear parabolic equations including the normalized and variational $p$-Laplace equations and their further generalizations, as well as regularizing approximations for degenerate $p$-Laplace equations ($1\leq p\leq\infty$). Below we summarize our major convergence rate results obtained for these equations. See Section 4 for precise statements, detailed analysis, concrete examples, and further discussions.
@@
Note that the Hölder regularity of solutions assumed in the results above is not the focus of this paper. We refer interested readers to [16, 22, 1, 20, 29], among many others, for related regularity estimates about these parabolic equations.
@@ Acknowledgments
The authors would like to thank Hiroyoshi Mitake and Félix del Teso for helpful discussions. The work of QL was supported by JSPS Grant-in-Aid for Scientific Research, No. 22K03396.
@@ 2. Preliminaries
Note that if $A$ is continuous in ${\mathbb{R}}^{n}$, then the class $\mathcal{F}$} contains all functions $f(r)=r^{k}$ with $k>2$, and the corresponding $\mathcal{F}$}-solutions are known to coincide with the standard viscosity solutions in this non-singular case; see for example [19, Proposition 2.28] for the proof of this fact. Hence, Definition 2.2 can be viewed as a generalized notion of viscosity solutions that applies also to parabolic equations with singularities at vanishing gradients.
@@ 3. Quantitative stability
Step 2. Application of the Crandall-Ishii Lemma. Assume thriving for a contradiction that both $(x_{\delta},t_{\delta}),(y_{\delta},s_{\delta})\in\Omega_{T}$. Then we can use the $(n+1)$-dimensional elliptic-type Crandall-Ishii Lemma [11, Theorem 3.2] (or [34, Lemma 3.5]) to obtain $(\tau,\xi,X)\in\overline{P}^{2,+}u_{\varepsilon}(x_{\delta},t_{\delta})$, $(\sigma,\eta,Y)\in\overline{P}^{2,-}u(y_{\delta},s_{\delta})$ for each $\mu>0$ such that
@@ 4.4. Regularization for generalized $p$-Laplace equations
Our result above, in the case of $p^{\prime}=2$, is consistent with [33, Theorem 1] for Lipschitz solutions to the level set mean curvature flow equation, which corresponds to our setting with $p^{\prime}=2$, $p=1$, $\theta=1$ and $g_{\varepsilon}=g_{0}$. Despite the different settings of boundary conditions, Theorem 4.7 verifies the same convergence rate $O(\varepsilon^{\nu})$ with any $\nu\in(0,1/2)$, as obtained in [33]. Our result additionally provides the exponent bound $\theta/2$ when the solutions $u_{\varepsilon}$ are merely $\theta$-Hölder continuous in space.

### v2

@@ 1.1. Motivation
In this paper we study the stability for viscosity solutions of a class of quasilinear parabolic partial differential equations, which take the form
@@
are given functions satisfying appropriate assumptions to be introduced later. Here, ${\mathcal{S}}^{n}_{+}$ represents the set of all nonnegative symmetric $n\times n$ matrices. A key feature of this general setting is that the elliptic operator may exhibit singularities at vanishing gradients.
@@
One of the main questions addressed in this paper concerns the stability of $p$-Laplace type equations under perturbations of the parameter $p$. For a fixed parabolic boundary value $g$, it is known that, as $p\to q$ for some $q\in(1,\infty)$, the unique solution $u_{p}$ of the Cauchy-Dirichlet problem (1.2) or (1.3) converges uniformly to the solution $u_{q}$ of the corresponding problem with parameter $q$. In the framework of viscosity solutions, consult [35, Theorem 6.1] and [20, Theorem 2.2.1] for such qualitative stability results on the variational and normalized $p$-Laplace equations of parabolic type. See also [29, 33] for related analysis with different approaches. However, these works do not provide an explicit rate of convergence. Our goal is therefore to further quantify the uniform convergence and to discuss the optimality of the resulting estimates.
On the other hand, when studying the solutions $u$ of singular parabolic equations such as (1.2) and (1.3), it is common to approximate the solutions of regularized equations of the form
@@
where $\varepsilon>0$ is small. Here, the choices $p^{\prime}=2$ and $p^{\prime}=p$ yield regularized approximations of (1.2) and (1.3), respectively. As $\varepsilon\to 0$, the comparison principle for the limiting equation yields the uniform convergence of the solution $u_{\varepsilon}$ of (1.5) to the solution $u$ of the corresponding singular equation. It is then natural to ask about the rate of this convergence in applications. The case with $p^{\prime}=2$ and $p=1$ corresponds to the level-set mean curvature flow equation, for which such a regularization was introduced in [18]. The convergence rate in this case was essentially established by Deckelnick [14], and later proved by Mitake [34] using a different method based on comparison arguments. Addressing convergence rates for approximations as in (1.5) is another main objective of the present paper.
The method employed in this paper builds on the arguments developed by Mitake in [34], as well as earlier works by Cockburn, Gripenberg and Londen [10] and Jakobsen and Karlsen [25], which established general continuous dependence estimates for viscosity solutions of nonlinear parabolic equations. Further related results include extensions to nonlocal PDEs in [26] and to general Neumann-type boundary value problems in [24]. While it may be possible to adapt these methods to study quantitative stability for $p$-Laplace type equations in regimes without vanishing-gradient singularities, precise estimates do not appear to have been explicitly given, to the best of our knowledge. Also, the singular case, such as (1.3) with $1<p<2$, seems unexplored and constitutes one of the focuses of the present work.
A recent paper of Bungert [5] studies the convergence rate of $p$-harmonic functions to infinity-harmonic functions, using partly the variational structure of the $p$-Laplace operator. In contrast, based on the viscosity solution theory, we here provide a purely PDE approach to show the convergence rate for a general class of related parabolic stability problems. Our method further develops the techniques in [10, 25, 34] and extends them to treat convergence rates arising from more general stability and regularization problems.
@@ 1.2. Main results
These assumptions guarantee the existence and uniqueness of continuous $\mathcal{F}$}-solutions to the associated Cauchy-Dirichlet problem; see [20]. The existence of $k>2$ required in (A1) is an assumption related to possible singularities of the equation at vanishing gradients. A typical example of singular operators we will consider in our later study is the case of the stability of the $p$-Laplacian for $p$ near a fixed $q\in(1,2)$, where $A_{\varepsilon}$ is taken to be
@@
While the Hölder regularity of solutions is not the primary focus of this paper, our method relies on the equi-Hölder continuity of $\mathcal{F}$}-solutions up to the boundary. In general, the validity of such estimates may depend on the equation structure, boundary data, and the geometry of $\Omega$. In the case of $p$-Laplace type parabolic equations, we refer the reader to [16, 23, 1, 21] for related interior regularity results, and to [17, 30] for boundary estimates. Note that both bodies of literature are extensive, and we cite only a selection here. We also remark that the regularity theory for the general equation (1.1) remains only partially understood.
@@ 1.3. Applications
Our general framework in Theorem 1.1 yields stability estimates for a broad class of quasilinear parabolic equations including the normalized (1.2) $p$-Laplace equation and variational $p$-Laplace equation (1.3) and their further generalizations, as well as regularizing approximations for degenerate $p$-Laplace equations ($1\leq p\leq\infty$). Below we summarize our major convergence rate results obtained for these equations. See Section 4 for precise statements, detailed analysis, concrete examples, and further discussions.
@@ Acknowledgments
The authors would like to thank Hiroyoshi Mitake, Félix del Teso and Denis Brazke for helpful discussions. The work of QL was supported by JSPS Grant-in-Aid for Scientific Research, No. 22K03396.
@@ 2. Preliminaries
Note that if $A$ is continuous in ${\mathbb{R}}^{n}$, then the class $\mathcal{F}$} contains all functions $f(r)=r^{k}$ with $k>2$, and the corresponding $\mathcal{F}$}-solutions are known to coincide with the standard viscosity solutions in this non-singular case; see for example [20, Proposition 2.28] for the proof of this fact. Hence, Definition 2.2 can be viewed as a generalized notion of viscosity solutions that applies also to parabolic equations with singularities at vanishing gradients.
@@ 3. Quantitative stability
Step 2. Application of the Crandall-Ishii Lemma. Assume thriving for a contradiction that both $(x_{\delta},t_{\delta}),(y_{\delta},s_{\delta})\in\Omega_{T}$. Then we can use the $(n+1)$-dimensional elliptic-type Crandall-Ishii Lemma [11, Theorem 3.2] (or [35, Lemma 3.5]) to obtain $(\tau,\xi,X)\in\overline{P}^{2,+}u_{\varepsilon}(x_{\delta},t_{\delta})$, $(\sigma,\eta,Y)\in\overline{P}^{2,-}u(y_{\delta},s_{\delta})$ for each $\mu>0$ such that
@@ 4.4. Regularization for generalized $p$-Laplace equations
Our result above, in the case of $p^{\prime}=2$, is consistent with [34, Theorem 1] for Lipschitz solutions to the level set mean curvature flow equation, which corresponds to our setting with $p^{\prime}=2$, $p=1$, $\theta=1$ and $g_{\varepsilon}=g_{0}$. Despite the different settings of boundary conditions, Theorem 4.7 verifies the same convergence rate $O(\varepsilon^{\nu})$ with any $\nu\in(0,1/2)$, as obtained in [34]. Our result additionally provides the exponent bound $\theta/2$ when the solutions $u_{\varepsilon}$ are merely $\theta$-Hölder continuous in space.